please tell me if the body initial velocity is zero then s=ut + 1/2at^2 i find that time =under root 2s/a but we know that v=s/t and a=v/t then from these two equation time=under root s/a please explainme why the result is different in both the cases

Grade:11

1 Answers

animesh sinha
13 years ago

here, there is a problem with concepts.

s=ut+1/2att

u=0

so,t=under root 2s/a

a=v/t

till here everything is fine.

v=s/t

here, v=average velocity

this v is not equal to v used above

v used above=instantaneous velocity

Hope u got ur answer